# Multiscale contact mechanics for elastoplastic contacts

### v1

@@ Abstract
Understanding contact between rough surfaces undergoing plastic deformation is crucial in many applications. We test Persson’s multiscale contact mechanics theory for elastoplastic solids, assuming a constant penetration hardness. Using a numerical model based on the boundary element method, we simulate the contact between a rough, rigid surface and an elastic-perfectly plastic half-space with a flat surface. The theory’s predictions for elastic, plastic, and total contact area agree quantitatively with the numerical results. The simulations also support the boundary conditions assumed in the theory, namely that the stress probability vanishes at both zero and yield stress. These findings reinforce the validity of the theory for systems with constant hardness.
Corresponding author: B.N.J. Persson,
email: b.persson@fz-juelich.de
1  Introduction
@@
One analytical approach to the elastoplastic contact between solids is the Persson contact mechanics theory. This theory, in its simplest form, assumes that the solids yield plastically when the local stress reaches the penetration hardness $\sigma_{\ rmP}$. The penetration hardness is determined in indentation experiments, where a very hard object (often diamond) of well-defined shape (usually a ball or pyramid) is pressed into contact with the solid. The ratio between the loading force and the area of the (plastically deformed) indentation is defined as $\sigma_{\rm P}$. The elastoplastic contact mechanics theory has been applied to the leakage of metallic seals[1] and to the leakage of syringes with Teflon-laminated rubber stoppers[2]. In a recent study, Lambert and Brodsky[3] have applied the theory (with a length-scale dependent penetration hardness) to the contact between the surfaces in earthquake faults.
The Persson contact mechanics theory is based on studying the probability distribution $P(\sigma,\zeta)$ of contact stress $\sigma$ as the magnification $\zeta$ increases. This function obeys a diffusion-like equation where time is replaced by magnification $\zeta$ and the spatial coordinate by the stress $\sigma$, and where the diffusivity depends on the surface roughness power spectrum and on the elastic properties of the solids. Solving this equation requires boundary conditions. Since there can be no negative stress at the interface (assuming no adhesion) $P(\sigma,\zeta)=0$ for $\sigma<0$. In fact, one can show that for elastic contact $P(\sigma,\zeta)\rightarrow 0$ as $\sigma\rightarrow 0^{+}$. For elastoplastic contact, the same condition is assumed to hold, with the additional assumption that $P(\sigma,\zeta)\rightarrow 0$ as $\sigma\rightarrow\sigma_{\rm P}^{-}$, where $\sigma_{\rm P}$ is the penetration hardness—defined as the ratio between the indentation force and the projected contact area—and is typically related to the yield stress, $\sigma_{\rm Y}$, by $\sigma_{\rm P}\approx 3\sigma_{\rm Y}$.
The original derivation of the stress probability distribution in the asperity contact regions for elastoplastic contact resulted in an infinite series. In a recent study, Xu et al.[4] have derived a simpler analytical expression for the stress probability distribution. This study also claims to have proved the correctness of the boundary conditions $P(\sigma,\zeta)\rightarrow 0$ as $\sigma\rightarrow 0^{+}$ and $\sigma\rightarrow\sigma_{\rm P}^{-}$. However, the proof is not clear to us, it appears to make use of the solution obtained, assuming the correctness of the boundary conditions.
@@
The aim of this study is to perform a rigorous test of the Persson elastoplastic theory in the simplest case of a constant penetration hardness. We will compare the theory predictions to exact boundary element calculations, which are possible for systems with roughness over a small length scale region. We will show that the theory predicts accurately both the elastic and plastic contact area, and also supports the assumption that $P(0,\zeta)=P(\sigma_{\rm P},\zeta)=0$ are the correct boundary conditions to use when solving the stress probability diffusion equation.
@@
2  Multiscale contact mechanics theory
We study the interface between an elastoplastic solid with a perfectly flat surface and a rigid body with random roughness at a magnification $\zeta$. At this magnification, we only observe the roughness with wavenumber components below $\zeta q_{0}$, where $q_{0}$ is the wavenumber of the smallest roughness component. We assume that the solids have nominally flat surfaces and are squeezed together with a nominally uniform stress. The nominal contact region is denoted by $A_{0}$. Because of the surface roughness, the stress acting at the interface is highly non-uniform. In the Persson contact mechanics theory for elastic solids, the probability distribution of contact stress equals:
@@ 2  Multiscale contact mechanics theory
where $C(q)$ is the surface roughness power spectrum. When the relative contact area $A/A_{0}<<1$, as in the applications below, $S(q)\approx\gamma$ (see Ref. [7]). The elastic energy reduction factor $\gamma$ has been determined originally by comparing the average surface separation as a function of squeezing pressure with numerical simulation results, and has been found to be $\gamma\approx 0.5$. Here we will use $\gamma=0.536$.
When the stress in the asperity contact region becomes high enough, plastic flow occurs. In the simplest model, it is assumed that a material deforms as a linear elastic solid, until the stress reaches a critical level, known as the plastic yield stress, at which it flows without strain hardening, as shown in Fig. 1. The yield stress in elongation is denoted by $\sigma_{\rm Y}$. In indentation experiments, where a sharp tip or a sphere is pushed against a flat solid surface, the penetration hardness $\sigma_{\rm P}$ is defined as the ratio between the normal force and the projected (on the surface plane) area of the plastically deformed indentation. Typically, $\sigma_{\rm P}\approx 3\sigma_{\rm Y}$. We note that the yield stress of materials often depends on the length scale (or magnification), which in principle can be included in the formalism we use [8, 3].
The influence of plastic flow on the contact mechanics can be taken into account in the Persson contact mechanics approach by replacing the boundary condition $P(\infty,\zeta)=0$ with the condition that there is no stress at the interface above the penetration hardness, i.e., $P(\sigma,\zeta)=0$ for $\sigma>\sigma_{\rm P}$. Thus, the maximum stress at the interface is equal to the penetration hardness $\sigma_{\rm P}$. This approach is based on the simplest elastoplastic description, where only elastic deformation occurs for $\sigma<\sigma_{\rm P}$, while for $\sigma=\sigma_{\rm P}$, the material flows without work-hardening so that the maximal stress equals $\sigma_{\rm P}$ (see Fig. 1).
The pressure probability distribution for this case in the region where elastic deformation has occurred is given by[9]:
$P(\sigma,\zeta)=\frac{2}{\sigma_{\rm P}}\sum_{n=1}^{\infty}\,{\rm sin}(s_{n}\sigma_{0})\,{\rm sin}(s_{n}\sigma)\,e^{-s_{n}^{2}G(\zeta)},$ (3)
where $s_{n}=n\pi/\sigma_{\rm P}$. If $A_{\rm el}(\zeta)/A_{0}$ is the fraction of the nominal area where the solids are in elastic contact when the interface is observed at the magnification $\zeta$, then
@@
$=\frac{2}{\pi}\sum_{n=1}^{\infty}\,\frac{1}{n}\left[1-(-1)^{n}\right]{\rm sin}(s_{n}\sigma_{0})\,e^{-s_{n}^{2}G(\zeta)}$ (4)
Note that, as $\sigma_{\rm P}\rightarrow\infty$, (4) reduces to (1). The fraction of the nominal area where the solids are in plastic contact when the interface is observed at the magnification $\zeta$ is
$\frac{A_{\rm pl}(\zeta)}{A_{0}}=\frac{\sigma_{0}}{\sigma_{\rm P}}+\frac{2}{\pi}\sum_{n=1}^{\infty}\,\frac{(-1)^{n}}{n}{\rm sin}(s_{n}\sigma_{0})\,e^{-s_{n}^{2}G(\zeta)}.$ (5)
3  Boundary element method
The numerical model employed in this work simulates two-dimensional frictionless contact between a rigid, randomly rough surface and an elastic-perfectly plastic half-space, using a spectral boundary element method grounded in the framework introduced by Almqvist et al. [10]. Surface roughness is represented as a periodic height field, and deformations are computed in the frequency space, via convolution with the Green’s function of an elastic half-space. Plastic deformation is introduced through a local yield pressure: wherever the contact pressure exceeds this threshold, the node is treated as plastically loaded and constrained to maintain contact without bearing further load. The solver employs relaxation-based iteration with residual convergence to enforce force balance and has been furthered and applied in several other works, e.g., [11, 12, 13, 14, 15, 16, 17].
@@ 3  Boundary element method
Figure 3: The calculated (a) in-contact height topography and (b) contact stress distributions. The results are obtained using the boundary element (BE) method for an elastoplastic model with Young’s modulus $E=250\ {\rm GPa}$, Poisson ratio $\nu=0$, and the yield stress (or penetration hardness $\sigma_{\rm P}=150\ {\rm GPa}$. The applied squeezing pressure is $\sigma_{0}=11.25{\rm GPa}$. The calculated relative elastic contact area $A_{\rm el}/A_{0}=0.0542$ (yellow) and the plastic contact area $A_{\rm pl}/A_{0}=0.0568$ (black). The total contact area $A/A_{0}=(A_{\rm el}+A_{\rm pl})/A_{0}=0.1109$.
@@
Figure 4: The elastic and plastic contact area as a function of the magnification, as predicted by the analytical theory. The true contact areas are obtained at the highest magnification and are $A_{\rm el}/A_{0}=0.0533$ and $A_{\rm pl}/A_{0}=0.0572$. The total contact area $A/A_{0}=(A_{\rm el}+A_{\rm pl})/A_{0}=0.1105$. For the same applied pressure and the same material parameters as in Fig. 3. The elastic energy reduction factor $\gamma=0.536$.
@@
Figure 5: The dotted red line is the probability distribution of contact stresses for elastic contact with the same parameters as in Fig. 3 except with $\sigma_{\rm P}=\infty$. The red solid line is the probability distribution of elastic stress from the analytical theory. The green and blue lines are the probability distribution of elastic stress from the boundary element calculations for the 2048-8 and 4096-16 systems, respectively. In the calculations of the solid lines, we have used the same applied pressure and the same material parameters as in Fig. 3.
4  Comparison of theory with numerical results
@@
We have studied the contact mechanics of an elastoplastic solid with a flat surface being squeezed against a rigid substrate with a randomly rough surface. The rough surfaces have self-affine fractal roughness in a wavenumber interval $1\times 10^{5}\ {\rm m}^{-1}<q<64\times 10^{5}\ {\rm m}^{-1}$, with the fractal dimension $D_{\rm f}=2.2$ (corresponding to the Hurst exponent $H=0.8$). We have added a short roll-off region for $2.5\times 10^{4}\ {\rm m}^{-1}<q<1\times 10^{5}\ {\rm m}^{-1}$ as this generate some self averaging, and in addition, most surfaces of engineering interest have a roll-off region in the power spectra (since most surfaces are designed to be smooth at the macroscopic length scale). When comparing numerical simulations with continuum mechanics theories, it is crucial to have enough grid points within the shortest roughness wavelength. This can be obtained by using surfaces which are generated from power spectra that have a large enough cut-off region for large wavenumber, as illustrated in Fig. 2 for a surface where the “width” of the cut-off region is $16$. The total “length” of the power spectrum is $q_{1}/q_{0}=4096$, so we denote the surface as $4096-16$. The power spectrum shown in Fig. 2 has the rms-slope $\xi=1$. We will consider surfaces with power spectra identical to the one shown in Fig. 2, except for the differing lengths of the large wavenumber cut-off regions. All these surfaces have the rms-slope $\xi\approx 1$ since the cut-off region contributes negligibly to the rms slope.
@@
We have generated randomly rough surfaces $z=h(x,y)$ with the power spectra given above by adding plane waves with random phases:
$h({\bf x})=\sum_{\rm q}B({\bf q})e^{i[{\bf q}\cdot{\bf x}+\phi_{\bf q}]}$ (7)
where $\phi_{\bf q}$ are random numbers uniformly distributed between 00 and $2\pi$. In (7)
$B({\bf q})=\frac{2\pi}{L}[C({\bf q})]^{1/2},$ (8)
@@
Fig. 3(a) shows the calculated height topography when the elastic block is squeezed against a surface with the power spectrum 4096-16. The results are obtained using the boundary element method (BEM) for an elastoplastic model with the parameters given above. Fig. 3(b) shows the stress acting in the area of real contact. From the stress distribution, for the 8192-32 system, one can obtain the relative elastic contact area $A_{\rm el}/A_{0}=0.0533$, and the plastic contact area $A_{\rm pl}/A_{0}=0.0572$. The total contact area $A/A_{0}=(A_{\rm el}+A_{\rm pl})/A_{0}=0.1105$. We will compare the numbers to the theoretical prediction.
Fig. 4 shows the elastic and plastic contact area as a function of the magnification as predicted by the analytical theory (4) and (5). The true contact areas are obtained at the highest magnification and are $A_{\rm el}/A_{0}=0.0526$ and $A_{\rm pl}/A_{0}=0.0572$. The total contact area $A/A_{0}=(A_{\rm el}+A_{\rm pl})/A_{0}=0.1099$. These numbers are in nearly perfect agreement with the numerical simulation results, as shown in Table 1.
Table 1: Comparison of fractional elastic ($A_{\text{el}}/A_{0}$), plastic ($A_{\text{pl}}/A_{0}$), and total contact area ($A/A_{0}$) for surfaces with different cut-off lengths.
Cut-off length $A_{\text{el}}/A_{0}$ $A_{\text{pl}}/A_{0}$ $A/A_{0}$
BEM 8 0.0576 0.0551 0.1126
16 0.0542 0.0568 0.1109
32 0.0533 0.0572 0.1105
Theory 32 0.0526 0.0572 0.1099
Fig. 4 shows that at low magnification, where only the long-wavelength roughness can be observed, it appears as if the contact is purely elastic. Only at high enough magnification can one observe a plastically deformed contact area. Thus, the theory predicts the length scale at which plastic flow becomes important, providing very useful information in many applications.

### v3

@@ Abstract
Understanding contact between rough surfaces undergoing plastic deformation is crucial in many applications. We test Persson’s multiscale contact mechanics theory for elastoplastic solids, assuming a constant penetration hardness. Using a numerical model based on the boundary element method, we simulate the contact between a flat rigid surface and an elastic-perfectly plastic half-space with a randomly rough surface. The theory’s predictions for elastic, plastic, and total contact area agree quantitatively with the numerical results. The simulations also support the boundary conditions assumed in the theory, namely that the stress probability vanishes at both zero and yield stress. These findings reinforce the validity of the theory for systems with constant hardness.
I Introduction
@@
One analytical approach to the elastoplastic contact between solids is the Persson contact mechanics theory. This theory in its simplest form assumes that the solids yield plastically when the local stress reaches the penetration hardness $\sigma_{\rm P}$. The penetration hardness is determined in indentation experiments, where a very hard object (often diamond) of well-defined shape (usually a ball or pyramid) is pressed into contact with the solid. The ratio between the loading force and the area of the (plastically deformed) indentation is defined as $\sigma_{\rm P}$. The elastoplastic contact mechanics theory has been applied to the leakage of metallic seals [1] and to the leakage of syringes with Teflon laminated rubber stoppers [2]. In a recent study, Lambert and Brodsky [3] have applied the theory (with a length-scale dependent penetration hardness) to the contact between the surfaces in earthquake faults.
Consider the contact between two elastoplastic solid bodies with nominal flat surfaces with roughness. When the interface is observed at the magnification $\zeta$ only the roughness components with wavelength $\lambda>L/\zeta$, can be observed, where where $L$ is the linear size of the surface. We define the wavenumber $q=2\pi/\lambda$ and the reference wavenumber $q_{0}=2\pi/L$.
The Persson contact mechanics theory is based on studying the probability distribution $P(\sigma,\zeta)$ of contact stress $\sigma$ as the magnification $\zeta$ increases. This function obeys a diffusion-like equation where time is replaced by magnification $\zeta$ and the spatial coordinate by the stress $\sigma$, and where the diffusivity depends on the surface roughness power spectrum and on the elastic properties of the solids. Solving this equation requires boundary conditions. Since there can be no negative stress at the interface (assuming no adhesion) $P(\sigma,\zeta)=0$ for $\sigma<0$. In fact, one can show that for elastic contact $P(\sigma,\zeta)\rightarrow 0$ as $\sigma\rightarrow 0^{+}$. For elastoplastic contact, the same condition is assumed to hold, with the additional assumption that $P(\sigma,\zeta)\rightarrow 0$ as $\sigma\rightarrow\sigma_{\rm P}^{-}$, where $\sigma_{\rm P}$ is the penetration hardness. Here we will show, by comparing the analytical theory with exact numerical simulation results, that both boundary conditions hold for elastoplastic contact.
The original derivation of the stress probability distribution in the asperity contact regions for elastoplastic contact resulted in an infinite series [4]. In a recent study, Xu et al. [5] have derived a simpler analytical expression for the stress probability distribution.
@@
The aim of this study is to perform a rigorous test of the Persson elastoplastic theory in the simplest case of a constant penetration hardness. We compare the theory predictions to exact boundary element calculations, which are possible for systems with roughness over a small length scale region. We show that the theory predicts accurately both the elastic and plastic contact area, and also support the assumption that $P(0,\zeta)=P(\sigma_{\rm P},\zeta)=0$ are the correct boundary conditions to use when solving the stress probability diffusion equation.
@@
II Multiscale contact mechanics theory
We study the interface between an elastoplastic solid with a randomly rough surface [8], and a rigid body with a perfectly flat surface at a magnification $\zeta$. At this magnification, we only observe the roughness with wavenumber components below $\zeta q_{0}$, where $q_{0}$ is the wavenumber of the smallest roughness component. We assume that the solids have nominally flat surfaces and are squeezed together with a nominally uniform stress. The nominal contact area is denoted by $A_{0}$. Because of the surface roughness, the stress acting at the interface is highly non-uniform. In the Persson contact mechanics theory for elastic solids, the probability distribution of contact stress equals:
@@ II Multiscale contact mechanics theory
$G(\zeta)=\frac{\pi}{4}(E^{*})^{2}\int_{q_{0}}^{\zeta q_{0}}dq\ q^{3}C(q)S(q),$ (2)
where $E^{*}=E/(1-\nu^{2})$ is an effective elastic modulus and $C(q)$ the surface roughness power spectrum.
@@
When the stress in the asperity contact region becomes high enough, plastic flow occurs. In the simplest model, it is assumed that a material deforms as a linear elastic solid until the stress reaches a critical level, the so-called plastic yield stress, where it flows without strain hardening, see Fig. 1. The yield stress in elongation is denoted by $\sigma_{\rm Y}$. In indentation experiments, where a sharp tip or a sphere is pushed against a flat solid surface, the penetration hardness $\sigma_{\rm P}$ is defined as the ratio between the normal force and the projected (on the surface plane) area of the plastically deformed indentation. As shown by Tabor[10], typically $\sigma_{\rm P}\approx 3\sigma_{\rm Y}$. We note that the yield stress of materials often depends on the length scale (or magnification), which in principle can be included in the formalism we use [11, 3].
The influence of plastic flow on the contact mechanics is taken into account in the Persson contact mechanics approach by replacing the boundary condition $P(\infty,\zeta)=0$ with $P(\sigma_{\rm Y},\zeta)=0$. Note that although there can be no stress larger than the penetration hardness, i.e., $P(\sigma,\zeta)=0$ for $\sigma>\sigma_{\rm P}$, this does not by itself guarantee that $P(\sigma,\zeta)\rightarrow 0$ as $\sigma\rightarrow\sigma_{\rm P}^{-}$. The same is true as $\sigma\rightarrow 0^{-}$ so the boundary conditions $P(\sigma_{\rm Y},\zeta)=0$ and $P(0,\zeta)=0$ are non-trivial. This approach is based on the simplest elastoplastic description, where only elastic deformation occurs for $\sigma<\sigma_{\rm P}$, while for $\sigma=\sigma_{\rm P}$, the material flows without work-hardening so that the maximal stress equals $\sigma_{\rm P}$ (see Fig. 1).
The pressure probability distribution in the elastoplastic theory, in the region where elastic deformation has occurred, is given by [4]:
@@
where $s_{n}=n\pi/\sigma_{\rm P}$. Note that, as $\sigma_{\rm P}\rightarrow\infty$, (3) reduces to (1). If $A_{\rm el}(\zeta)/A_{0}$ is the fraction of the nominal area where the solids are in elastic contact when the interface is observed at the magnification $\zeta$, then
$\frac{A_{\rm el}(\zeta)}{A_{0}}=\int_{0}^{\sigma_{\rm P}}d\sigma\ P(\sigma,\zeta)$
[…]
The fraction of the nominal area where the solids are in plastic contact when the interface is observed at the magnification $\zeta$ is
[…]
III Boundary element method
The numerical model employed in this work simulates two-dimensional frictionless contact between an elastic-perfectly plastic half-space with a randomly rough surface and a rigid body with a flat surface, using a spectral boundary element method grounded in the framework introduced by Almqvist et al. [12]. Surface roughness is represented as a periodic height field, and deformations are computed in the frequency space, via convolution with the Green’s function of an elastic half-space. Plastic deformation is introduced through a local yield pressure: wherever the contact pressure exceeds this threshold, the node is treated as plastically loaded and constrained to maintain contact without bearing further load. The solver employs relaxation-based iteration with residual convergence to enforce force balance and has been furthered and applied in several other works, e.g., [13, 14, 15, 16, 17, 18, 19].
@@ III Boundary element method
Figure 3: The calculated (a) in-contact topography squeezed against a rigid flat surface (transparent), and (b) contact stress distributions. The results are obtained using the boundary element method (BEM) for an elastoplastic model with Young’s modulus E=250GPaE=250\ $\mathrm${GPa}, Poisson ratio $\nu=0$, and the yield stress (or penetration hardness) $\sigma_{\rm P}=150\ \mathrm{GPa}$. The applied squeezing pressure is $\sigma_{0}=12.50\ \mathrm{GPa}$. The calculated relative elastic contact area is $A_{\rm el}/A_{0}=0.0533$ (yellow), plastic contact area $A_{\rm pl}/A_{0}=0.0572$ (black), and total contact area $A/A_{0}=0.1105$.
@@
Figure 4: The elastic and plastic contact area as a function of the magnification (log-log scale), as predicted by the analytical theory. The true contact areas are obtained at the highest magnification and are $A_{\rm pl}/A_{0}=0.0572$ and $A_{\rm el}/A_{0}=0.0526$. The total contact area $A/A_{0}=(A_{\rm el}+A_{\rm pl})/A_{0}=0.1099$. For the same applied pressure and the same material parameters as in Fig. 3. The elastic energy reduction factor $\gamma=0.536$.
@@
Figure 5: The black dotted line is the probability distribution of contact stresses for elastic contact with the same parameters as in Fig. 3 except with $\sigma_{\rm P}=\infty$. The block solid line is the probability distribution of elastic stress from the analytical theory. The red, green and blue lines are the probability distribution of elastic stress from the boundary element calculations for the 2048-8, 4096-16 and 8192-32 systems, respectively. Not shown in the figure are the Dirac delta functions $\sim\delta(\sigma)$ and $\sim\delta(\sigma-\sigma_{\rm P})$, corresponding to the non-contact area and plastically deformed contact area, respectively. In the calculations of the solid lines, we have used the same applied pressure and the same material parameters as in Fig. 3.
IV Comparison of theory with numerical results
@@
We have studied the contact mechanics of an elastoplastic solid with a randomly rough surface being squeezed against a rigid substrate with flat surface. The rough surfaces have self-affine fractal roughness in a wavenumber interval $1\times 10^{5}\ {\rm m}^{-1}<q<64\times 10^{5}\ {\rm m}^{-1}$, with the fractal dimension $D_{\rm f}=2.2$ (corresponding to the Hurst exponent $H=0.8$). We have added a short roll-off region for $2.5\times 10^{4}\ {\rm m}^{-1}<q<1\times 10^{5}\ {\rm m}^{-1}$ as this generate some self averaging, and in addition, most surfaces of engineering interest have a roll-off region in the power spectra (since most surfaces are designed to be smooth at the macroscopic length scale) [20]. When comparing numerical simulations with continuum mechanics theories, it is crucial to have enough grid points within the shortest roughness wavelength. This can be obtained by using surfaces which are generated from power spectra that have a large enough cut-off region for large wavenumber, as illustrated in Fig. 2 for a surface where the “width” of the cut-off region is $16$. The total “length” of the power spectrum is $q_{1}/q_{0}=4096$, so we denote the surface as 4096-16. The power spectrum shown in Fig. 2 has the rms-slope $\xi=1$. We consider surfaces with power spectra identical to the one shown in Fig. 2, except for the differing lengths of the large wavenumber cut-off regions. All these surfaces have the rms-slope $\xi\approx 1$ since the cut-off region contributes negligibly to the rms slope.
@@
We have generated randomly rough surfaces $z=h(x,y)$ with the power spectra given above by adding plane waves with random phases (see Appendix A in Ref. [21]):
[…]
where $\phi_{\bf q}$ are random numbers uniformly distributed between 0 and $2\pi$. In (7)
[…]
@@
Fig. 3(a) shows the calculated height topography when the elastic block is squeezed against a surface with the power spectrum 4096-16. The results are obtained using the boundary element method (BEM) for an elastoplastic model with the parameters given above. Fig. 3(b) shows the stress acting in the area of real contact. From the stress distribution, for the 8192-32 system one can obtain the relative elastic contact area $A_{\rm el}/A_{0}=0.0533$, and the plastic contact area $A_{\rm pl}/A_{0}=0.0572$. The total contact area $A/A_{0}=(A_{\rm el}+A_{\rm pl})/A_{0}=0.1105$. We compare these values with the theoretical predictions.
Fig. 4 shows the elastic and plastic contact area as a function of the magnification as predicted by the analytical theory (4) and (5). The true contact areas are obtained at the highest magnification and are $A_{\rm el}/A_{0}=0.0526$ and $A_{\rm pl}/A_{0}=0.0572$. The total contact area $A/A_{0}=(A_{\rm el}+A_{\rm pl})/A_{0}=0.1099$. These numbers are in nearly perfect agreement with the numerical simulation results, see Table 1. In the table we also show the relative contact area $A/A_{0}$ for elastic contact, i.e., when $\sigma_{\rm Y}=\infty$, but with all the other material parameters the same as for the elastoplastic contact. The elastic contact mechanics case has been studied in great detail elsewhere [22, 23] and will not be discussed further here.
elastoplastic elastic
Cut-off length $A_{\mathrm{el}}/A_{0}$ $_{\mathrm{pl}}/A_{0} AA$$_{0} AA$$A_{0}$
BEM 8 0.0576 0.0551 0.1126 0.1028
16 0.0542 0.0568 0.1109 0.1003
32 0.0533 0.0572 0.1105 0.1002
Theory 0.0526 0.0572 0.1099 0.1087
Table 1: Comparison of fractional elastic ($A_{\mathrm{el}}/A_{0}$), plastic ($A_{\mathrm{pl}}/A_{0}$), and total contact area ($A/A_{0}$) for elastoplastic contact for surfaces with different cut-off lengths. The last column shows the same for elastic contact.
Fig. 4 shows that at low magnification, where only the long-wavelength roughness can be observed, it appears as if the contact is purely elastic. Only at high enough magnification can one observe a plastically deformed contact area. Thus, the theory predicts the length scale at which plastic flow becomes important, providing very useful information in many applications. Note also that when the magnification becomes so large that the $\zeta q_{0}$ enter the cut-off region in Fig. 2 the contact area does not change any more as the magnification increases. This is due to the (nearly) vanishing amplitude of these short wavelength roughness components.